jericakayla:

Which fraction is equivalent to negative 4 over 6? 6 over negative 4 negative 6 over 4 4 over negative 6 4 over 6

dude:

Any idea?

jericakayla:

no thats why im asking I think is b or d

jericakayla:

im not sure though

dude:

b and d look right, are the actual fractions different or was it written?

jericakayla:

what do u mean?

jericakayla:

I know its not a

dude:

Which fraction is equivalent to negative 4 over 6? A. \(\frac{6}{-4}\) B. \(\frac{-6}{4}\) C. \(\frac{4}{-6}\) D. \(\frac46\)

dude:

Is it written like that?

jericakayla:

yea

jericakayla:

well its the negative sign with the 4 then 6

dude:

Well the question asks \(\frac{-4}{6}\) So the fraction has to remain with a 4 on top and 6 on bottom but the signs can change as long as only 1 of them has it

jericakayla:

so c?

dude:

Yep

jericakayla:

k thx

dude:

Sure